OpenStudy (anonymous):

A major stadium has 45,000 seats. If baseball have bought 5/8 of the seats for the next game, how many seats are avalible ?

OpenStudy (agent47):

45,000-(45,000*5/8)

OpenStudy (anonymous):

thatwas wrong , But THANK YOU ( :

OpenStudy (mayankdevnani):

set up a proportion if 5/8 of the 45,000 seats are sold,then 5/8 = x/45,000 then cross multiply or use cross products, so 5(45,000) = 8x 225,000 = 8x ----------- --- (divide by 8) 8 8 28,125 = x <== that is 5/8 of the seats, or how many are sold. To find out how many are available, subtract the number of seats already sold (28,125) from the number of seats there are in the stadium (45,000) so, 45,000 - 28,125 = 16875 seats are still available.

OpenStudy (mayankdevnani):

@breannaw

OpenStudy (anonymous):

thank uu

OpenStudy (mayankdevnani):

most welcome.....